Nancy Cotton bought 400 shares of NeTalk for $15 per share. One year later, Nancy sold the stock for $21per share, just after she received a $0.90 cash dividend from the company.

a) What is dollar return earned by Nancy for the year?

b) What is the rate return earned by Nancy?

c) separate the rate of return computed in part (b) into the dividend yield and the capital gains yield. In other words, compute the dividend yield and the capital gains yield that Nancy earned by holding NeTalk for one year.
Answer #1

a)

dollar return = number of shares * (sell price - buy price + dividend)

= 400 * (21 -15 + 0.9)

= 2760

b)

rate of return = dollar return per share/buy price

= (21 -15 + 0.9)/15

= 0.46

= 46%

c)

dividend yield = 0.9/15 = 0.06 = 6%

capital gains yield = (21-15)/15 = 0.4 = 40%
Answer #2

a- 465

b-98%

c-74
Answer #3

No of shares = 400 shares 

Purchase Value of share = $15 per share. 

Selling value of the share = $21per share

Dividend = $0.90 


a) What is dollar return earned by Nancy for the year?

Dollar Return = Number of shares * (Selling Price - Purchase price + dividend)

= 400 * ($21 -$15 + $0.9)

= $2760


b) What is the rate return earned by Nancy?

Rate of return = Dollar return per share/ Purchase price

Rate of return = ($21 -$15 + $0.9)/$15

Rate of return = $6.9/$15

Rate of return = 0.46 or 46%


c) Separate the rate of return computed in part (b) into the dividend yield and the capital gains yield. In other words, compute the dividend yield and the capital gains yield that Nancy earned by holding NeTalk for one year.

Dividend yield = Declared dividend/Purchase price

Dividend yield = 0.9/15

Dividend yield = 0.06 or 6%

Capital gains yield = (Selling Price - Purchase price)/ Purchase price

Capital gains yield = (21-15)/15 

Capital gains yield = 0.4 or 40%


answered by: Puneet Goel